### v1

# A probabilistic foundation model for crystal structure denoising, phase classification, and order parameters

Abstract
Atomistic simulations generate large volumes of noisy structural data, but extracting phase labels, order parameters (OPs), and defect information in a way that is universal, robust, and interpretable remains challenging. Existing tools such as PTM and CNA are restricted to a small set of hand-crafted lattices (e.g. FCC/BCC/HCP), degrade under strong thermal disorder or defects, and produce hard, template-based labels without per-atom probability or confidence scores. Here we introduce a log-probability foundation model that unifies denoising, phase classification, and OP extraction within a single probabilistic framework. We reuse the MACE-MP foundation interatomic potential on crystal structures mapped to AFLOW prototypes, training it to predict per-atom, per-phase logits $l$ and to aggregate them into a global log-density $\log\hat{P}_{\theta}(\bm{r})$ whose gradient defines a conservative score field. Denoising corresponds to gradient ascent on this learned log-density, phase labels follow from $\arg\max_{c}l_{ac}$, and the $l$ values act as continuous, defect-sensitive and interpretable OPs quantifying the Euclidean distance to ideal phases. We demonstrate universality across hundreds of prototypes, robustness under strong thermal and defect-induced disorder, and accurate treatment of complex systems such as ice polymorphs, ice–water interfaces, and shock-compressed Ti.
Introduction
Atomistic simulations are central tools for studying solid–solid and solid–liquid phase transitions, defect formation, and microstructural evolution in materials [belonoshko2006melting, zepeda2017probing, shibuta2017heterogeneity, Zepeda-Ruiz2017N]. Advances in first-principles calculations, machine-learning interatomic potentials (MLIPs), and high-performance computing now enable routine multi-million atom simulations over long timescales. However, extracting physical insight from such datasets still hinges on two challenging analysis tasks: (i) assigning crystalline phase labels to individual atoms, and (ii) defining continuous order parameters (OPs) that quantify the degree of structural order and track phase transformations. For realistic, thermally perturbed configurations with defects, surfaces, grain boundaries, or partial melting, systematic and universally applicable tools for these tasks are still lacking.
Significant progress has been made on crystal structure classification for ideal or weakly perturbed unit cells. The Curtarolo group, for example, has curated the AFLOW Encyclopedia of structural prototypes [Mehl2017CMS-prototypeP1, Hicks2018CMS-prototypeP2, Hicks2021CMS-prototypeP3, Eckert2024CMS-prototypeP4] and developed tools such as XtalFinder [Hicks2021nCM-XtalFinder], which efficiently match relaxed primitive cells to known prototypes. For large-scale atomistic configurations, a range of local structural descriptors is widely used, including common neighbor analysis (CNA) [Honeycutt1987JPC-CNA], bond-orientational OPs [Steinhardt1983PRB, Lechner2008JCP], centrosymmetry analysis [Kelchner1998PRB-CSA], and polyhedral template matching (PTM) [larsen2016robust]. These methods are highly effective for a handful of well-studied lattices such as BCC, FCC, and HCP, and have become standard in analysis packages like OVITO [Stukowski2012MSMSE-review]. Yet, they typically rely on hand-crafted geometric thresholds and domain-specific heuristics, limiting their transferability to complex or less common prototypes. Under strong thermal distortions, disorder, or coexistence of multiple phases, they often mislabel atoms or return ambiguous classifications [Hsu2024nCM-denoiser].
Continuous OPs provide complementary scalar measures of structural order. Classical examples include Steinhardt-type bond-order parameters and related metrics for liquid–solid transitions [Steinhardt1983PRB, Lechner2008JCP]. However, unlike the AFLOW prototype catalog for crystal structures, no analogous, systematic “encyclopedia” of OPs exists. Instead, OPs are typically designed on a case-by-case basis, tailored to particular polymorphs or specific transitions (e.g., FCC/BCC/HCP). This lack of a general framework hinders automated analysis of large, heterogeneous datasets and complicates thermodynamic characterization of complex phase behavior.
Machine learning (ML) offers an attractive path toward more general structure characterization. Early work combined symmetry-invariant descriptors (e.g., SOAP, bispectrum) with neural networks to classify crystal structures or detect phase transitions [Ziletti2018NC, Geiger2013JCP, Defever2019CS, Fulford2019JCIM, Kim2020PCCP, Swanson2020SM, Doi2020JCP, Becker2021PRE, Leitherer2021NC, Chung2022PRM, Hernandes2022JPCM, Chapman2022nCM, Chapman2023NC, Aroboto2023APL, Moradzadeh2023JPCC]. In our previous works [Hsu2024nCM-denoiser, Sun2024JCIM-ice], we adapted the score-based diffusion models [Sohl-Dickstein2015-DPM, Ho2020-DDPM, Song2019NeurIPS-Generative] from generative AI to atomistic systems, treating thermal noise removal as a statistical inference problem. A machine-learned denoiser model approximates the non-conservative score (nominally the gradient of a log-density) of ideal crystalline configurations and uses it to iteratively remove thermal perturbations from noisy structures [Hsu2024nCM-denoiser, Sun2024JCIM-ice]. Coupled with conventional classifiers (e.g., CNA and PTM), this two-stage pipeline achieved near-perfect phase classification for a few familiar phases up to the melting point, while preserving physically meaningful disorder such as defects.
Despite these successes, existing scientific ML approaches still exhibit several severe limitations for broad applications. First, denoising and classification have typically been viewed and designed as separate tasks: a denoising model is trained for the purpose of either noise removal [Hsu2024nCM-denoiser, Sun2024JCIM-ice] or featurization/pretraining [Zaidi2023ICLR, New2024-crystaldenoising, Shen2024-denoising-pre-training], without explicit knowledge of classification objectives, and a downstream classifier operates only on the cleaned structures. This separation complicates training and may discard subtle structural information useful for discrimination between closely related phases (e.g., HCP vs. $\omega$). Second, most methods focus on producing discrete labels, with limited use of per-atom probabilities or confidence scores to expose ambiguity. This is particularly problematic near phase boundaries, in highly disordered regions, or for structures outside the training distribution. Third, many models are system-specific, specialized to a small set of phases or chemistries, and not ostensibly generalizable to arbitrary crystalline prototypes.
In face of these limitations, an ideal framework for structural analysis should therefore satisfy three criteria simultaneously. First, it should be universal, operating across a wide range of crystal prototypes and chemistries rather than being restricted to a few hand-tuned lattices such as FCC/BCC/HCP. Second, it must be robust to realistic perturbations such as thermal noise, defects, interfaces, and out-of-equilibrium configurations that are ubiquitous in large-scale simulations and experimental reconstructions. Third, it should offer interpretable outputs. Existing symmetry-based, fingerprinting, and task-specific ML methods typically satisfy at most one or two of these requirements.
Energy-based models (EBMs) provide a natural and unifying statistical perspective for addressing these goals [LeCun2006-Energy, Grathwohl2020ICLR, Li2023-DiffusionClassifier, Chen2023-RobustClassifier]. In an EBM, a scalar “energy” function $E(\bm{r})$ defines a (usually unnormalized) probability distribution $P(\bm{r})\propto\exp[-E(\bm{r})]$, such that the gradient or score $\partial_{\bm{r}}\log\!P$ drives sampling or denoising. This viewpoint suggests that a single model could, in principle, assign probabilities to multiple crystal phases, yield a score field for denoising, and define OPs through its scalar outputs. However, existing applications of EBMs and diffusion models to atomistic systems have not yet fully realized this joint denoising–classification–OP potential, and have largely focused on either generative sampling or noise removal.
In this work, we take a step toward such a unified framework by introducing a probabilistic model that simultaneously denoises atomic configurations, classifies crystalline phases, and provides continuous OPs. The model predicts per-atom, per-phase logits (unnormalized scores) $l_{ac}$ for each atom $a$ and candidate phase $c$. Aggregating these across atoms via a log-sum-exp yields a total machine-learned log-density $\log\!\hat{P}_{\theta}(\bm{r})$, whose gradient defines a conservative score field $\bm{s}(\bm{r})=\partial_{\bm{r}}\log\!\hat{P}_{\theta}(\bm{r})$ for denoising, while the per-phase logits serve as physically motivated OPs measuring similarity to each class $c$. Phase labels are obtained directly by selecting the class $c$ with the largest $l_{ac}$, and ambiguous regions can be identified through low or mixed $l_{ac}$ values. Training follows the paradigm of MLIPs, combining a denoising score-matching loss [Vincent2011scorematching] (analogous to force matching) and a cross-entropy classification loss on the logits $l_{ac}$ (loosely relatable to energy matching). In contrast, our previous denoiser model directly predicts a non-conservative score field $\hat{\bm{s}}_{\theta}(\bm{r})$, similar to direct-force predictions of some force fields, with no explicit conservative log-probability structure [Hsu2024nCM-denoiser, Sun2024JCIM-ice].
Here we introduce a log-probability foundation model for crystalline materials: a single equivariant neural network trained across hundreds of AFLOW prototypes and thousands of elemental and binary structures, designed to serve as a reusable backbone for diverse downstream structure-analysis tasks. In this domain-specific sense, “foundation” refers to broad structural coverage and transferability across phases, rather than to hyperscale web- or text-scale training typical of language models.
Practically, we instantiate this idea by reusing the MACE-MP foundation model [Batatia2022mace, Batatia2023foundation] via transfer learning with fixed featurization on a curated subset of Materials Project structures [jain2013commentary] mapped to AFLOW prototypes [hicks2021aflow]. The training dataset includes elemental, binary, and ternary crystals, augmented with random elastic strains and Gaussian positional noise to mimic realistic thermal and mechanical perturbations. The resulting log-probability foundation model achieves near-perfect prototype classification and sub-Å denoising errors across hundreds of crystal types, maintains high accuracy on benchmark datasets of thermally perturbed configurations, and generalizes to challenging out-of-distribution scenarios such as shock-compressed Ti with coexisting BCC, HCP, FCC, and $\omega$ phases, as well as water–ice interfaces with mixed solid–liquid regions. These results demonstrate that a log-probability foundation model can provide a general, data-efficient route to automated structure recognition and probabilistic OPs for noisy atomistic configurations, while meeting the universality, robustness, and probabilistic interpretability criteria outlined above.
Figure 1: Overview of the log-probability ($\log\!P$) foundation model. Training uses ideal crystalline structures mapped to AFLOW prototypes, with two coupled objectives: (i) predicting per-atom, per-class logits $l_{ac}$ guided by crystal class labels, and (ii) learning the conservative score field $\partial_{\bm{r}}\log\!\hat{P}_{\theta}(\bm{r})$ of the aggregated log-density $\log\!P$ from randomly displaced structures. At inference time, the same model can be used to iteratively denoise noisy configurations, assign phase labels from $\operatorname*{arg\,max}_{c}l_{ac}$, and evaluate per-atom $l_{ac}$ fields as continuous, phase-resolved OPs.
Results
Before discussing applications, we briefly summarize how the $\log\!P$ model defines OPs. The network predicts per-atom, per-phase logits (unnormalized scores) $l_{ac}$, from which we construct a global log-density
$\log\hat{P}_{\theta}(\bm{r})=\sum_{a}\log\sum_{c}\exp\left(\hat{l}_{\theta;ac}(\bm{r})\right).$ (1)
Here and throughout, we refer to $l_{ac}$ as logits and reserve “(log-)probability” for the aggregated quantity $\log\hat{P}_{\theta}(\bm{r})$ (or its normalized softmax over classes when needed). The per-phase logits $l_{ac}$ act as continuous, phase-resolved OPs that quantify how similar each atom $a$ is to prototype $c$, and $\operatorname*{arg\,max}_{c}l_{ac}$ provides categorical classification. The gradient
$\hat{\bm{s}}(\bm{r})=\partial_{\bm{r}}\log\hat{P}_{\theta}(\bm{r})$ (2)
defines a conservative score field used for denoising. In what follows we refer to these per-atom $l$ values as probabilistic OPs. This architecture, which predicts per-atom scalar outputs (logits for phases, cf. energies in an MLIP), allows us to leverage well-established training pipelines of MLIPs using derivative (score or force) matching. The details can be found in the Method section.
Foundation model performance on large crystalline dataset
We first assess the performance of the log-probability foundation model on the curated Materials Project dataset described in the Methods. The model is trained jointly for denoising and crystalline prototype classification: given a noisy atomic configuration, it predicts per-atom, per-class logits (unnormalized scores) and a conservative score field whose gradient is used to iteratively refine atomic positions (Figure 1). This shared log-probability landscape underlies both the denoising dynamics and the final phase assignments.
Table 1 summarizes performance across representative subsets of the dataset, including ice polymorphs, elemental, binary, and ternary compounds spanning hundreds of AFLOW prototypes. The model achieves near-perfect classification accuracy and sub-Å denoising errors across all tested systems. For the combined elemental+binary set (7,746 structures, 403 structure types), the foundation model reaches classification accuracies above 99.9% on clean inputs and maintains similarly high accuracy on Gaussian-perturbed structures with a noise standard deviation of 0.15 Å, while keeping the denoising RMSE below 0.002 Å. Notably, the chemistry-agnostic elemental model—which shares a single ML representation across all elements—still attains $\sim$96% accuracy, indicating that the ML descriptors capture robust geometric information even without explicit chemical labels.
The chemistry-agnostic model is especially important for extending the approach to high-entropy alloys and other compositionally complex systems, where many elements can share the same lattice sites. In such settings, template-based methods and chemistry-specific models must be retrained or reparameterized for each composition, whereas the geometry-only probabilistic model can directly recognize the underlying prototype regardless of the particular elemental labels.
Material # structures # prototypes # atom types Class. acc. at step 8 (clean / perturbed 0.15 Å) RMSE (Å) Class. acc. at step 0
Ice 7 7 2 1.0000 / 1.0000 0.0191 / 0.0191 1.0000 / 1.0000
Elemental structures 238 33 72 0.9961 / 1.0000 0.0002 / 0.0013 0.9961 / 0.9961
Binary structures 7488 363 75 0.9988 / 0.9983 0.0013 / 0.0019 0.9991 / 0.9991
Ternary structures 14848 373 84 0.9977 / 0.9937 0.0019 / 0.0020 0.9981 / 0.9972
Elemental + binary structures 7746 403 75 0.9993 / 0.9991 0.0009 / 0.0011 0.9994 / 0.9990
Elemental chemistry-agnostic 238 33 1 0.9625 / 0.9595 0.0054 / 0.0091 0.9628 / 0.9699
Table 1: Performance of the log-probability foundation model on the curated Materials Project dataset. The model jointly learns to denoise atomic coordinates and classify crystal prototypes across ice polymorphs, elemental, binary, and ternary compounds. For each dataset, we report the number of structures, prototypes, and atom types, together with classification accuracy on clean and perturbed inputs (Gaussian noise up to 0.15 Å), denoising RMSE, and accuracy at step 0 (before any denoising steps are applied). The model achieves near-perfect accuracy and sub-Å denoising errors across all crystalline systems.
We leveraged the strong expressive power of the MACE-MP foundation model by reusing its featurization layers and adding a new trainable decoder that predicts per-atom logits and the aggregated $\log\!P$. This transfer-learning setup significantly accelerates convergence compared to training from scratch (see Supplementary Fig. S.1 and Methods).
Multi-phase denoising and classification in ice polymorphs
Figure 2: Per-phase logit distributions $l_{ac}$ for 7 ice polymorphs (Ic, Ih, II, III, VI, VII, and sI). Each row corresponds to a given true ice phase and each column to a predicted structural class. Within each panel, the light curves show the initial perturbed structures, intermediate curves show partially denoised configurations (after 3 of 8 denoising steps), and the darkest curves show the fully denoised structures. Diagonal panels (true class = predicted class) develop sharp, high-$l$ peaks as denoising proceeds, indicating confident and self-consistent phase recognition, while off-diagonal panels remain suppressed at low $l$.
As a first multi-phase test, we apply the foundation model to 7 ordered ice polymorphs (Ic, Ih, II, III, VI, VII, and sI), which provide a familiar but nontrivial benchmark with distinct hydrogen-bonding networks and local environments. The model is trained jointly on all 7 phases and evaluated on Gaussian-perturbed structures with noise amplitudes up to $\sigma_{\max}=0.15$ Å, using the same denoising protocol as for the crystalline solids. The Gaussian displacements mimic thermal-like positional fluctuations around the ideal lattice sites.
Figure 2 shows the distributions of per-atom logits $l_{ac}$ for each input phase (rows) and predicted structural class (columns), at different stages of the denoising process. Light-colored curves correspond to the initial perturbed configurations, intermediate curves correspond to partially denoised structures (e.g., after 3 out of 8 denoising steps), and the darkest curves represent the fully denoised outputs. Along the diagonal panels–where the predicted class matches the true phase–the $l_{ac}$ distributions develop pronounced peaks at high values as denoising proceeds, indicating confident and self-consistent classification. Off-diagonal panels remain narrowly peaked at lower $l_{ac}$, reflecting smaller weights assigned to incorrect phases.
Quantitatively, the model achieves perfect classification accuracy (1.000) for all seven ice phases, both for clean inputs and for perturbed structures with $\sigma_{\max}=0.15$ Å, while maintaining denoising RMSEs on the order of $2\times 10^{-2}$ Å (Table 1). These results demonstrate that a single probabilistic model can robustly distinguish multiple hydrogen-bonded phases even under substantial thermal-like perturbations. They also illustrate how the per-phase logits $l_{ac}$ naturally act as continuous OPs: each phase is associated with a distinct, well-separated logit distribution that sharpens under denoising, providing a scalar measure of structural similarity suitable for tracking phase identity and transformation pathways. In contrast, a separate, second-stage descriptor-based classifier was needed to supplement the non-conservative denoiser in Ref. [Sun2024JCIM-ice].
Interpretable OPs and continuous transformation paths
We next examine how the foundation model behaves on familiar close-packed structures and along continuous deformation paths between them. This serves both as a sanity check that the machine-learned $\log\!P$ landscape respects well-known crystallographic relationships and as a quantitative test of the physical interpretability of the logit-based OPs.
Figure 3 focuses on a silver structure in the hexagonal A_hP2_194 (space group 194) prototype. Panel 3a illustrates the effect of coupling or decoupling denoising from classification for a strongly perturbed configuration with Gaussian positional noise of 0.4 Å. When the model is applied in a purely classification mode (no denoising), the heavily distorted local environments cause the structure to be misclassified into a variety of competing prototypes. When denoising and classification are performed jointly through the shared $\log\!P$ landscape, the atomic positions are iteratively refined toward high-$\log\!P$ regions, and the model converges back to the correct A_hP2_194 class even from this strongly perturbed starting point.
Panel 3b directly probes the approximate quadratic relation between the logits and the displacement from the ideal reference structure derived in the Methods section. It plots the mean logit for the correct phase, $\langle l(\mathrm{A\_hP2\_194})\rangle$, versus the mean squared displacement per atom, $\langle|\Delta\bm{r}|^{2}\rangle$, along the denoising trajectory for multiple initial noise levels. The data align closely with a linear trend (shown by a regression line), consistent with the denoising score-matching setup (see Method)
$l\approx\mathrm{const}-\lVert\Delta\bm{r}\rVert^{2}/(2\sigma^{2})=\mathrm{const}-\lVert\bm{r}-\bm{R}_{0}\rVert^{2}/(2\sigma^{2})$ (3)
relative to the correct ideal phase $\bm{R}_{0}$. The plot provides explicit evidence that the learned logits inherit a direct physical meaning as distance-like OPs measuring proximity to the corresponding ideal prototype.
Panel 3c shows how the per-phase logit distributions evolve during denoising for 3 closely related close-packed prototypes, $l(\mathrm{A\_hP3\_166})$, $l(\mathrm{A\_hP4\_194})$, and $l(\mathrm{A\_hP2\_194})$, at denoising steps 0, 3, and 7. At step 0, the noisy structure exhibits broad, partially overlapping logit distributions, and the correct A_hP2_194 class is not clearly dominant. After a few denoising steps, the logit distribution for A_hP2_194 sharpens and shifts to higher values, while the competing phases are suppressed and pushed toward lower logits. By step 7, the correct class forms a well-separated high-$l$ peak, and the impostor phases remain narrowly distributed at low $l$. This illustrates how the logit-based OPs act as continuous, phase-resolved measures of structural similarity that naturally become more decisive as the structure is projected onto the learned high-probability manifold.
Finally, panel 3d summarizes the net effect on predictive performance by plotting the classification accuracy and final denoising RMSE as functions of the initial Gaussian noise standard deviation. The model maintains 100% classification accuracy for perturbations up to 0.5 Å, with small denoising errors, and both metrics degrade beyond this point as the structures melt and no longer correspond to well-defined crystalline phases. Together, panels (a)–(d) show that the $\log\!P$ model not only stabilizes classification through denoising but also yields logit-based OPs that vary smoothly and approximately quadratically with the squared distance to the underlying prototype, in line with the intended probabilistic interpretation.
Figure 3: Interplay between denoising, classification, and logit-based OPs for noisy Ag in the A_hP2_194 prototype. (a) Example with strong Gaussian noise ($\sigma=0.4$ Å). In classification-only mode (no denoising), the structure is misclassified into several competing AFLOW prototypes. When denoising and classification are coupled through the log-probability foundation model, the atomic positions are iteratively refined toward high-$\log\!P$ regions and the correct A_hP2_194 label is recovered for all atoms. (b) Mean logit for the correct prototype, $\langle l(\mathrm{A\_hP2\_194})\rangle$, versus mean squared displacement per atom, $\langle|\Delta\bm{r}|^{2}\rangle$, for a range of initial noise levels and denoising steps. The approximately linear trends (dashed regression lines) are consistent with the local Gaussian model in which $l$ is proportional to the negative squared distance to the ideal structure, providing a direct physical interpretation of the logit-based OP. (c) Evolution of per-phase logit distributions for 3 closely related prototypes, $l(\mathrm{A\_hP3\_166})$, $l(\mathrm{A\_hP4\_194})$, and $l(\mathrm{A\_hP2\_194})$, at denoising steps 0, 3, and 7. As denoising proceeds, the logit distribution for the correct A_hP2_194 phase sharpens and shifts to higher values, while competing phases are suppressed, illustrating how the logit-based OPs become more decisive as the structure is projected onto the high-probability manifold. (d) Classification accuracy and final denoising RMSE as a function of the initial Gaussian noise standard deviation. The model maintains 100% accuracy up to $\sigma\approx 0.5$ Å, beyond which both accuracy and denoising quality degrade as the structures melt and no longer correspond to well-defined crystalline phases.
To probe whether the $\log\!P$ model captures smooth structural evolution between phases, we evaluate it along two standard transformation paths: the Bain path connecting BCC and FCC, and the Burgers path connecting HCP and BCC (Figure 4). Along each path, we generate a sequence of intermediate configurations with gradually changing lattice parameters and atomic positions. For each configuration, we evaluate the per-atom, per-phase logits $l_{ac}$ and aggregate them into prototype-resolved OPs.
The resulting profiles of these per-phase logits and their differences demonstrate their usefulness as continuous, physically interpretable OPs. Along the Bain path, the BCC logit-based OP starts high in the initial BCC-like region and decreases monotonically as the structure is distorted toward FCC, while the FCC logit-based OP rises in a complementary fashion and dominates near the FCC endpoint. Similarly, along the Burgers path, the HCP logit-based OP decreases as the structure is driven toward BCC, whose logit-based OP increases and eventually becomes dominant. This smooth exchange of OP weight between competing phases indicates that the model does not treat prototypes as discrete, disconnected categories, but instead learns a continuous OP landscape over configuration space that tracks gradual structural transformations. Having emerged naturally from the crystalline structures alone, without requiring access to the underlying physics (e.g. an energy landscape) or detailed chemistry, these OPs can be defined in a consistent and universal way with a direct physical meaning related to the squared distance to the corresponding ideal phases.
Figure 4: Evolution of logit-based OPs along continuous transformation paths. The foundation model is evaluated along (left) the Bain path connecting BCC and FCC and (right) the Burgers path connecting HCP and BCC, using sequences of intermediate structures. For each configuration, per-phase logits for the competing structures are evaluated and aggregated into prototype-resolved OPs. Along the Bain path, $l(\mathrm{BCC})$ decreases while $l(\mathrm{FCC})$ increases, crossing smoothly near the midpoint; along the Burgers path, $l(\mathrm{HCP})$ decays as $l(\mathrm{BCC})$ rises. The smooth exchange of logit-based OP weight between phases shows that the model captures continuous structural evolution rather than treating prototypes as isolated categories.
Robustness to thermal disorder and point defects
We next evaluate the robustness of the foundation model under realistic thermal disorder and local defects, where traditional geometric descriptors often struggle. For thermal effects, we use the DC3 database [chung2022data], which contains molecular dynamics (MD) snapshots of elemental and binary crystals equilibrated near their respective melting points. These configurations exhibit substantial vibrational amplitudes and partial loss of local order, providing a stringent test of structural recognition under near- and above-melting conditions.
Figure 5 compares the classification performance of the log-probability foundation model against two widely used template-based methods, PTM and CNA. For each DC3 system, we take the highest-temperature snapshot available and consider a sequence of configurations obtained by applying $k=0,\dots,8$ denoising steps with the foundation model (with $k=0$ corresponding to the original DC3 snapshot). At each step $k$, we evaluate the log-probability model, PTM, and CNA on the same coordinates, i.e. on the configuration after $k$ denoising steps. Across all tested systems, our model reaches higher accuracy with fewer denoising iterations than PTM or CNA, and in many cases achieves 100% accuracy by step 3. At intermediate steps, where the structures remain visibly noisy, the log-probability model already outperforms both baselines, indicating that accurate structural recognition emerges early in the denoising process. This behavior reflects a key difference in philosophy: PTM and CNA rely on fixed geometric thresholds tuned to specific lattices, whereas the foundation model learns to associate a broad distribution of local environments with each prototype through its probabilistic training.
Our foundation model is not always the most accurate. For example, in the HCP panels of Fig. 5, the few-step classification accuracy falls behind PTM. This is because the model also predicts logits $l$ for many closely related close-packed prototypes such as A_hP3_191 and A_cF4_225, and with zero or few denoising steps the significant thermal perturbations lead to misclassifications into those competing phases. This is the price to pay for a foundation model that encompasses a large set of competing structures. In contrast, PTM considers only two close-packed phases, FCC and HCP. More broadly, it is conceivable that for specific applications involving only a few candidate phases, our foundation model may be less accurate than specialized classical or ML classifiers at zero or very few denoising steps. In such cases, one may either perform more denoising iterations or fine-tune the model on the restricted phase set.
Figure 5: Classification accuracy on thermally perturbed structures from the DC3 database. Top panels: accuracy versus denoising step for representative elemental and binary systems at high temperatures above their melting points, comparing the log-probability foundation model to PTM and CNA. The foundation model reaches higher accuracy within fewer denoising iterations and often achieves 100% accuracy by step 3. Bottom panel: summary of classification accuracy at step 2 across all tested systems, showing consistently superior early-stage performance over PTM and CNA under strong thermal disorder. For each denoising step $k$, PTM and CNA are evaluated on the same coordinates as the log-probability model, i.e. on the configuration obtained after $k$ log-probability denoising steps (with $k=0$ corresponding to the original DC3 snapshot).
To probe sensitivity to point defects, we introduce vacancy-type defects into a BCC Fe supercell by randomly removing a small fraction of atoms (5 and 10 vacancies out of 432 atoms, respectively; Figure 6a and b). These missing atoms distort the local environments around the defect cores and frequently cause PTM to misclassify neighboring atoms as FCC or label them as “unknown” (left panels), reflecting the fragility of hard, template-based labels under local coordination changes. In contrast, the log-probability foundation model correctly assigns all atoms to the BCC prototype for both vacancy concentrations (middle panels), preserving the global phase identity. At the same time, the continuous BCC logit-based OP provides a natural defect-sensitive measure of local order: when atoms are colored by their BCC logit value $l_{a,\mathrm{BCC}}$ (right panels), the undisturbed crystal interior appears uniformly bright (high $l_{a,\mathrm{BCC}}$), while shells surrounding the vacancies show localized depressions in $l_{a,\mathrm{BCC}}$ (darker purple), indicating reduced confidence and stronger local disorder. Thus, the model simultaneously maintains robust global phase recognition and yields a smooth, quantitative measure of local deviations from ideal BCC order that discrete template matching cannot provide.
Figure 6: Classification and log-probability OP in defective BCC Fe. Two BCC Fe supercells with randomly removed atoms are used to create vacancy-type point defects: (a) 5 vacancies and (b) 10 vacancies out of 432 atoms. In each case, the left panel shows PTM classification, where atoms near the vacancy cores are frequently misclassified as FCC or labeled as “unknown,” reflecting the sensitivity of hard, template-based labels to local distortions. The middle panels show the log-probability foundation model classification via $\operatorname*{arg\,max}_{c}l_{ac}$, in which essentially all atoms are correctly identified as BCC. The right panels color atoms by $l_{a,\text{BCC}}$, with lighter purple indicating higher $l_{a,\text{BCC}}$ (bulk-like BCC environments) and darker purple indicating lower $l_{a,\text{BCC}}$ (locally perturbed environments near defects). The bulk interior remains uniformly light, while vacancy neighborhoods appear as darker halos, demonstrating that the model simultaneously recovers the correct global BCC prototype and provides a continuous, defect-sensitive OP that highlights local deviations from ideal crystalline order.
Generalization to diverse binary prototypes
While many structure-identification methods are tuned to a small set of familiar lattices (e.g., BCC, FCC, HCP), the AFLOW prototype library contains a much broader spectrum of low-symmetry and less common structures. To assess whether the foundation model extends beyond close-packed metals and simple oxides, we evaluate it on binary systems with multiple polymorphs and nontrivial AFLOW labels.
Figure 7 illustrates two representative examples. Panel 7a shows an AgO structure in the AB_mP8_14 prototype, starting from a perturbed configuration and followed through successive denoising steps. At early iterations (e.g., step 1), some atoms are transiently assigned to alternative prototype classes such as A2B3_oF40_43, AB4_cP40_205, or AB2_cP6_224, reflecting local environments that momentarily resemble competing motifs. As denoising proceeds, these inconsistencies vanish and the model converges to a self-consistent assignment in which all atoms are correctly classified as the target AB_mP8_14 prototype.
Figure 7: Generalization of the log-probability foundation model to diverse binary prototypes. (a) AgO in the AB_mP8_14 prototype: starting from a perturbed configuration, the model progressively denoises the structure while tracking per-atom prototype labels. At early steps, some atoms are transiently assigned to competing prototype classes (e.g., A2B3_oF40_43, AB4_cP40_205, AB2_cP6_224), but these inconsistencies vanish as denoising proceeds and all atoms converge to the correct AB_mP8_14 class. (b) 5 ZnO polymorphs: for each prototype (rows), the evolution of per-atom logit-based OP distributions across denoising steps (columns) shows sharpening, well-separated peaks for the true class and suppressed values for competing classes. These examples highlight that the approach is not limited to simple BCC/FCC/HCP lattices but extends to low-symmetry AFLOW prototypes in binary systems.
Panel 7b considers 5 distinct ZnO polymorphs. For each prototype (rows), we track the evolution of per-atom logit distributions across denoising steps (columns). The light curves correspond to the initial perturbed structures, intermediate curves show partially denoised states (step 3 of 8), and the darkest curves represent the fully denoised outputs. In all cases, the logit distributions for the true prototype sharpen into a dominant, well-separated peak, while the competing classes remain suppressed. Together, these examples demonstrate that the log-probability foundation model is not restricted to a small set of canonical lattices, but readily generalizes to diverse, low-symmetry AFLOW prototypes in binary systems.
Log-probability OPs in mixed solid–liquid water–ice systems
To probe the behavior of the log-probability foundation model in heterogeneous environments with coexisting ordered and disordered regions, we apply it to a water–ice interface featuring solid–liquid coexistence (Figure 8). The model is trained on 7 ordered ice polymorphs (Ic, Ih, II, III, VI, VII, and sI); liquid water is therefore out-of-distribution and expected to appear as low in all ice-related logits. The interface configuration is obtained from an equilibrated water–ice molecular dynamics simulation at 300 K and 1 kbar, so that thermal fluctuations naturally introduce positional disorder throughout the system. This setting is challenging for traditional local OPs because strong structural gradients and finite-temperature fluctuations blur the distinction between crystalline and liquid-like environments, particularly near the interface.
Figure 8: Probabilistic OPs in a mixed water–ice interface. (a) Application of the foundation model in classification-only mode (no denoising) yields a clear separation between crystalline ice and liquid water, but a small fraction of ice-like environments are spuriously assigned to competing polymorphs such as Ic. (b) When denoising is enabled during inference, the crystalline region relaxes toward a high-value manifold of the Ih logit-based OP (large $l_{a,\mathrm{Ih}}$), correcting these misclassifications between ice polymorphs, while the disordered liquid region remains diffuse and low in all ice logits. (c) The OVITO CHILL+ algorithm, based on geometric thresholds, fails to robustly identify the crystalline Ih region and assigns a mixture of ice- and hydrate-like labels, with substantial portions of the liquid misclassified as ordered, underscoring the improved robustness of the log-probability foundation model in heterogeneous, high-entropy environments.
When applied directly to the finite-temperature configuration without denoising, the model already captures the broad distinction between crystalline ice and liquid water: atoms in the ice slab carry high logits for Ih or related ice polymorphs, whereas atoms in the liquid region are low in all ice logits (Figure 8a). The main residual errors arise inside the crystalline region, where a small number of Ih-like environments are misclassified as Ic due to local perturbations that transiently make them resemble cubic-ice environments.
Enabling denoising during inference mainly improves this polymorph assignment rather than the basic solid–liquid separation. As atoms in the crystalline region are iteratively moved toward higher log-probability configurations, their Ih logits increase and spurious Ic assignments are removed, yielding a nearly uniform Ih phase in the solid region. At the same time, atoms in the liquid region remain low in all ice-related logits (Figure 8b). The resulting spatial distribution of the Ih logit-based OP therefore acts as a smooth probabilistic indicator of ice-like order that is robust to both thermal noise and polymorph confusion.
For comparison, the OVITO CHILL+ algorithm, which is commonly used to distinguish ice from liquid water based on geometric criteria, fails to reliably identify the crystalline Ih region in this configuration and assigns a noisy mixture of ice- and hydrate-type labels, with substantial parts of the liquid misclassified as ordered (Figure 8c). In contrast, the log-probability foundation model, trained only on ideal ice polymorphs and synthetically perturbed configurations, remains robust in this mixed-phase, non-periodic setting. This example highlights how per-atom logit-based OPs derived from the log-probability model can serve as probabilistic OPs for complex interfacial systems, with potential applications to solid–liquid coexistence, nucleation, and interfacial free-energy estimation [hoyt2001method].
Out-of-distribution shock-compressed Ti
As an extreme out-of-distribution test, we apply the log-probability foundation model to a large-scale simulation of shock-compressed Ti that exhibits severe deformation and complex phase coexistence (Figure 9). This configuration is not included in training and contains a heterogeneous mixture of close-packed (HCP-like) and $\omega$-like regions under highly nonequilibrium conditions, with additional highly strained environments that some classifiers label as BCC-like. This makes it a stringent benchmark for generalized structural recognition. An ablation study in which we retrain the model without elastic-strain augmentation shows that strongly shocked HCP Ti is then systematically misclassified as the rhombohedral AFLOW prototype A_hR3_166 (space group 166), confirming that strain augmentation is essential to avoid spuriously interpreting elastic distortions as phase changes (Supplementary Fig. S.2).
Figure 9: Shock-compressed Ti as an out-of-distribution test. (a) Classification using the full log-probability foundation model that includes chemical species. (b) Spatial map of the maximum per-atom logit over all phases from the full model, with vivid regions indicating atoms that strongly match a single prototype (high $\max_{c}l_{ac}$) and whitish regions indicating atoms whose environments are difficult to assign confidently to any single phase (low $\max_{c}l_{ac}$), e.g. near phase boundaries or in strongly distorted regions. (c) Classification using the chemistry-agnostic foundation model that ignores species and relies purely on geometry. (d) PTM classification baseline applied directly to the original shock-compressed configurations without log-probability denoising. Because PTM does not include an explicit $\omega$ prototype, structurally $\omega$-like regions are mapped onto HCP, BCC, or “unknown”, and large parts of the microstructure are either misclassified or left unassigned. Both log-probability foundation models recover HCP-like and $\omega$-like domains and resolve coexisting regions under strong strain and disorder, while PTM fails to reliably distinguish the $\omega$ phase and attributes many $\omega$-like atoms to BCC.
Template-based methods such as PTM are fundamentally constrained in this regime because their prototype sets typically do not include the $\omega$ phase. Under strong strain and disorder, PTM predominantly identifies the underlying close-packed lattice (usually HCP) or labels large regions as BCC or “unknown”, and atoms that are structurally $\omega$-like are necessarily mapped onto the nearest available templates or left unassigned (Figure 9d). As a result, the expected HCP $\rightarrow$ $\omega$ transformation under shock loading is not cleanly resolved in the PTM phase map, and some regions that are physically $\omega$-like are instead labeled as BCC.
In contrast, both the full-element and chemistry-agnostic versions of the log-probability foundation model recover a clear separation between HCP-like and $\omega$-like domains across the sample (Figure 9a,c). The model assigns elevated $\omega$ logit-based OP values in the high-pressure domains where the $\omega$ phase has formed, retains HCP-like logits where the original close-packed structure persists, and identifies a smaller fraction of atoms as BCC-like in highly sheared or interfacial regions. We emphasize that these BCC-like assignments should be interpreted as local environments with BCC-like coordination rather than as evidence for a thermodynamically stable BCC phase in this particular simulation.
Panel 9b illustrates how the logit field provides a confidence-based indicator for the classification. For each atom, we compute the maximum logit over all phases, $\max_{c}l_{ac}$; vivid regions correspond to atoms that strongly match a single prototype (high $\max_{c}l_{ac}$), while whitish regions have low $\max_{c}l_{ac}$ and are difficult to assign confidently to any phase. These low-$\max_{c}l_{ac}$ regions cluster around phase boundaries and highly distorted zones, naturally highlighting where the microstructure is structurally ambiguous or far from any ideal prototype.
Because PTM lacks an explicit $\omega$ prototype and the log-probability model includes one, the atoms labeled BCC-like by PTM and by the log-probability model do not coincide spatially. In PTM, many $\omega$-like atoms are projected onto the BCC template, whereas in the log-probability model most of those atoms are correctly assigned to $\omega$, with BCC-like labels confined to a smaller set of strongly deformed environments. Taken together, the phase maps and associated confidence fields provide a detailed view of the HCP $\rightarrow$ $\omega$ transformation, resolving where the $\omega$ phase nucleates and how domains grow and interact under shock loading. This level of spatially resolved phase information, which is difficult to obtain from PTM alone, is well-suited for subsequent analyses of $\omega$-phase nucleation and growth mechanisms in dynamically loaded Ti.
Discussion
Despite being trained exclusively on ordered crystalline structures mapped to AFLOW prototypes and augmented only with synthetic elastic and thermal perturbations, the log-probability foundation model demonstrates strong generalization across a wide range of structural variations, including thermal distortions, point defects, mixed solid–liquid interfaces, and shock-induced phase coexistence. By learning a global scalar log-density $\log\hat{P}_{\theta}(\bm{r})$ whose gradient defines the denoising direction, the model unifies three tasks that are typically treated separately: denoising perturbed configurations, assigning crystal phase labels, and providing continuous, physically interpretable OPs derived from the same per-phase logit landscape $l_{ac}$. Although our model is modest in size compared with hyperscale language or vision foundation models, it plays an analogous role within the atomistic domain by providing a reusable, phase-agnostic representation and log-probability decoder that transfer across hundreds of crystal prototypes and a range of downstream tasks.
This unified view leads to practical advantages over conventional symmetry-based and template-based approaches such as CNA or PTM. In noisy MD trajectories and high-temperature (e.g. DC3) snapshots at or above the nominal melting point, where hard geometric thresholds often fail, the log-probability model maintains high classification accuracy and can recover the correct prototype within a few denoising steps. Per-atom logits $l_{ac}$ have a simple and physical interpretation: the squared distance with respect to the ideal structure. They act as smooth OPs that track gradual structural transformations, as illustrated by the Bain and Burgers paths and by the spatial variation across a water–ice interface. In shock-compressed Ti, the model resolves coexisting HCP, BCC, and $\omega$ domains under strong strain and disorder, while template-based PTM, which lacks an explicit $\omega$ prototype, necessarily maps structurally $\omega$ regions onto HCP, BCC, or “unknown” labels. The resulting OP fields $l_{ac}$ and phase maps ($\operatorname*{arg\,max}_{c}l_{ac}$ with adjustable thresholds) provide a detailed description of phase coexistence and interfaces as well as physically easy-to-interpret OPs of similarity to structural prototypes, offering a natural starting point for quantitative analysis of $\omega$-phase nucleation and growth mechanisms in dynamically loaded Ti.
Our work is also closely related to earlier deep-learning approaches for crystal-structure classification, most notably the diffraction-image classifier of Ziletti et al. [Ziletti2018NC]. That study demonstrated that convolutional neural networks operating on 2D diffraction fingerprints can achieve nearly perfect classification of a small set of elemental crystal families and can remain robust under substantial disorder and defects. However, the classifier operates on reciprocal-space images and produces global class probabilities for a limited number of prototype classes. In contrast, the present log-probability foundation model works directly on real-space atomic graphs, scales to hundreds of AFLOW prototypes and thousands of elemental and binary structures, and outputs per-atom, per-phase $l$ values whose gradients define denoising displacements. This allows robust classification, denoising, and OP extraction to be handled within a single model, with spatial resolution sufficient to analyze interfaces, defects, and complex microstructures far beyond the scope of purely image-based classifiers.
A distinctive feature of the present approach is that it makes confidence and ambiguity in phase assignments directly visible. Regions that closely resemble a given prototype have large, sharply peaked $l$ for that class, whereas atoms near phase boundaries, defect cores, or strongly distorted environments exhibit reduced maxima or competing phase preferences. While this does not constitute a formal statistical uncertainty estimate in the sense of Bayesian or ensemble methods, it provides an intuitive, data-driven measure of how well each local environment matches the available prototypes. In practice, this graded view helps distinguish bulk-like regions from structurally atypical ones and complements hard categorical labels produced by existing tools.
Our current implementation has a practical computational limitation. The foundation model is built on the full MACE architecture used in MACE-MP, with relatively wide hidden representations and multiple interaction layers. While this choice is advantageous for accuracy and transferability, it also makes the model memory intensive. For very large atomistic configurations (e.g., shock simulations or large-scale MD snapshots with $10^{5}$ atoms), naive evaluation of the full model on a single GPU can lead to out-of-memory failures. In practice, this can be mitigated by domain decomposition and processing each subdomain separately, followed by stitching the predictions together. Another mitigation strategy is half-precision inference for large structures without discernible discrepancy compared to single or double-precision evaluations in our tests. Another limitation or burden of our foundation model is that it considers so many competing phases that its classification accuracy may be lower with zero or few denoising steps for “tricky” phases such as close-packed structures with different long-range stacking patterns. This can be easily solved with more denoising steps, or by focusing on outputs of a smaller pool of candidate structures. It is also possible that the frozen featurization layers of MACE-MP were relatively insensitive to subtle difference in long-range ordering, and therefore should be fine-tuned for improved classification accuracy.
Overall, this work establishes a unified, physically grounded paradigm for analyzing noisy atomic configurations. The log-probability foundation model does not only denoise structures; it provides a probabilistic framework that simultaneously explains, classifies, and quantifies structural order in crystalline materials, and that generalizes to challenging out-of-distribution cases such as high-temperature DC3 structures at or above the melting point and shock-compressed Ti. Because our model relies on no prior knowledge of specific crystalline phases other than the ideal structure, it can be straightforwardly generalized to quaternary and more complicated structures. Our probabilistic OPs, distinguished by their ease of development, universal applicability and direct physical meaning, will facilitate novel investigations in the modeling of phase transformations. Extending this framework to jointly model crystalline, liquid, and amorphous phases, to incorporate chemically disordered alloys, and to couple log-probability learning with generative sampling or automated prototype discovery are promising directions for future work. Such developments would further strengthen the role of log-probability foundation models as general tools for automated structure analysis, phase mapping, and data-driven thermodynamics in computational materials science.
Methods
Model architecture.
For clarity we reiterate the key definitions from equations (1,2). Given structure class $c$ (AFLOW prototype), the model predicts per-atom, per-class logits $\hat{l}_{\theta;ac}$ and aggregates them into a global log-probability and its associated conservative score field:
$\log\hat{P}_{\theta}(\bm{r})=\sum_{a}\log\sum_{c}\exp\left(\hat{l}_{\theta;ac}(\bm{r})\right),\ \hat{\bm{s}}(\bm{r})=\partial_{\bm{r}}\log\hat{P}_{\theta}(\bm{r})$
Phase labels are predicted by $\operatorname*{arg\,max}_{c}l_{ac}$ at inference time.
In practice, we instantiate this model by reusing the pretrained MACE-MP as a representation-learning backbone: the embedding and equivariant message-passing layers, jointly denoted as a featurization operator $\hat{F}$,
$\bm{z}_{a}=\hat{F}_{a}(\bm{r}),$ (4)
are kept frozen. A new trainable $C$-headed decoder $\hat{D}_{\theta}$, constructed similar to the original energy decoder, is added to predict per-atom logits from the learned latent representation $\bm{z}$:
${l}_{ac}=\hat{D}_{\theta,ac}(\bm{z}_{a}).$ (5)
This allows us to leverage the pretrained representation learned by MACE-MP while only retraining the final decoder layers for the $\log\!P$ objectives.
Dataset and augmentations.
To ensure consistency and relevance in structural representations, we curated a subset of the Materials Project [jain2013commentary] dataset by filtering entries to match crystal prototypes from the AFLOW Encyclopedia [hicks2021aflow]. Specifically, we include only Materials Project entries that (i) can be mapped to an AFLOW prototype and (ii) lie within 0.1 eV/atom above the convex hull, thereby focusing on experimentally plausible or metastable phases. The AFLOW Encyclopedia includes only prototypes observed in at least ten experimentally or computationally verified compounds, so this filtering step removes rare, idiosyncratic structures (e.g., CsMg149) that hinder generalization, and yields a dataset enriched in structurally meaningful crystalline motifs.
To account for physically realistic variations in lattice parameters, we applied a small random elastic deformation combining isotropic scaling and symmetric strain. For each structure, we first sampled an isotropic scale factor $s\sim\mathcal{U}[0.9,\,1.1]$ and then drew a strain tensor
$E_{ij}\sim\mathcal{U}(-\delta_{\text{strain}},\,\delta_{\text{strain}}),$ (6)
with $\delta_{\text{strain}}=0.05$. To avoid introducing spurious rigid-body rotations, we symmetrized the strain tensor as
$E\leftarrow\tfrac{1}{2}\,(E+E^{\mathsf{T}}),$ (7)
and formed the total deformation gradient
$T=s\,(I+E).$ (8)
The deformation $T$ was applied consistently to both the cell vectors and Cartesian atomic positions, followed by periodic wrapping of atoms back into the simulation cell. This augmentation exposes the model to moderate volumetric and shear strains while preserving the underlying prototype symmetry and periodicity.
Unless otherwise noted, we use $\sigma_{\max}=0.15$ Å as the maximum positional noise scale when constructing noisy configurations for score matching (see below). For each primitive cell, we build an approximately cubic supercell containing $\sim$210 atoms to provide sufficient local environments for graph-based learning.
Training objectives and optimization.
The total training loss combines a score-matching term and a classification term,
$\mathcal{L}$ $=\mathcal{L}_{\text{sm}}+w_{\text{cl}}\,\mathcal{L}_{\text{cl}},$ (9)
where $\mathcal{L}_{\text{cl}}$ encourages the logits $l_{ac}$ to match the known prototype label $c$ of the ideal structure $\bm{R}_{0}$, and $\mathcal{L}_{\text{sm}}$ enforces consistency with the Gaussian score.
During training, we construct noisy configurations by adding Gaussian noise to the ideal structure $\bm{R}_{0}=\{\bm{r}_{0a}\}_{a=1}^{N}$. For each structure, we first draw a noise amplitude
$\sigma_{n}\sim\mathcal{U}(0,\sigma_{\max}),$
then sample i.i.d. Gaussian noise
$\bm{\epsilon}\sim\mathcal{N}(0,\mathbf{I}),$
and define
$\tilde{\bm{R}}=\{\tilde{\bm{r}}_{a}\}_{a=1}^{N}=\bm{R}_{0}+\sigma_{n}\bm{\epsilon},\hskip 18.49988pt\Delta\bm{r}_{a}=\tilde{\bm{r}}_{a}-\bm{r}_{0a}.$
For classification, we use a per-atom cross-entropy loss,
$\mathcal{L}_{\text{cl}}=-\,\mathbb{E}_{\bm{R}_{0},c,\sigma_{n},\bm{\epsilon}}\left[\frac{1}{N}\sum_{a=1}^{N}\log\frac{\exp(l_{ac})}{\sum_{c^{\prime}}\exp(l_{ac^{\prime}})}\right].$ (10)
where $c$ is the AFLOW prototype label associated with $\bm{R}_{0}$.
[…]
$\mathcal{L}_{\text{sm}}=\mathbb{E}_{\bm{R}_{0},\sigma_{n},\bm{\epsilon}}\left[\frac{1}{N}\sum_{a=1}^{N}\left\|\hat{\bm{s}}_{a}(\tilde{\bm{R}})+\sigma_{n}^{-2}\,\Delta\bm{r}_{a}\right\|^{2}\right],$ (11)
[…]
Locally around a given prototype $c$, this Gaussian corruption model and score-matching objective encourage the network to approximate the conditional distribution of atomic displacements $\Delta\bm{r}_{a}=\bm{r}_{a}-\bm{R}_{0,a}^{(c)}$ as a Gaussian. In the simplest isotropic approximation,
$\log\hat{P}_{\theta,c}(\bm{r})\approx\mathrm{const}_{c}-\frac{1}{2\sigma_{c}^{2}}\sum_{a}\left\lVert\bm{r}_{a}-\bm{R}_{0,a}^{(c)}\right\rVert^{2},$ (12)
so that the per-atom logit for the correct class $c^{\star}$ behaves as
$l_{ac^{\star}}\approx\mathrm{const}_{c^{\star}}-\frac{1}{2\sigma_{c^{\star}}^{2}}\left\lVert\bm{r}_{a}-\bm{R}_{0,a}^{(c^{\star})}\right\rVert^{2}.$ (13)
Thus, up to an additive constant and a phase-dependent scale, $l_{ac^{\star}}$ is proportional to the negative squared distance between the current atomic position and the ideal reference position for phase $c^{\star}$. This provides a direct physical interpretation of the logit-based OPs as distance-like measures of similarity to each prototype. As shown for noisy Ag in the A_hP2_194 prototype in Fig. 3c, the learned logits indeed follow an approximately linear relation with both the mean squared displacement and the input noise variance, consistent with this local Gaussian picture.
Model parameters are optimized using AdamW with a learning rate of $1\times 10^{-3}$ and weight decay $1\times 10^{-4}$. For small datasets (fewer than 50 structures), training the $\log\!P$ decoder on top of a frozen MACE-MP backbone typically converges within about 2 hours on 4 nodes (4 GPUs per node). Across the full curated dataset, models initialized from MACE-MP foundation weights converge substantially faster than models trained from scratch (Supplementary Fig. S1), with the benefit most pronounced for larger MACE configurations (e.g., hidden irreps $128\times(0e+1o+2e)$ and 2 interaction layers). We attribute this to reusing the pretrained MACE-MP representation, including the scale and shift terms in the ScaleShiftBlock, which improves optimization stability and data efficiency.
Data availability
TBD
Code availability
The full code is available in the NPS repository at
Acknowledgment
This work was performed under the auspices of the U.S. Department of Energy by Lawrence Livermore National Laboratory under Contract DE-AC52-07NA27344. This work was funded by the Laboratory Directed Research and Development (LDRD) Program at LLNL under project tracking code 25-ERD-002 (HK, JK, FZ), 23-SI-006 (SH, VL). Computing support for this work came from the Lawrence Livermore National Laboratory Institutional Computing Grand Challenge program.
Author contributions
FZ developed and implemented the model and led the research. HK performed the computational experiments and data analysis. BS, SH, VL and JK contributed technical advice and data. JK and VL secured funding. HK and FZ wrote the paper with input from all authors.
Competing interests
The authors declare no competing interests.
Supplemental material
Appendix A Effect of foundation-model featurization on optimization.
To quantify the benefit of foundation-model pretraining, we compared two training protocols on the curated crystal-structure dataset. In the pretrained setup, we reuse the MACE-MP foundation model as a frozen equivariant featurizer and train only a newly added log-probability–based prototype-classification head on top of its representations. In the from-scratch setup, we use the same architecture but initialize all weights randomly and train end-to-end. As shown in Supplementary Fig. S.1, the model that reuses the MACE-MP featurization starts from lower initial classification and score-matching losses ($\mathcal{L}_{\text{cl}}$ and $\mathcal{L}_{\text{sm}}$) and converges more rapidly, reaching smaller final loss values in fewer epochs. This demonstrates that transferring a pretrained equivariant featurizer substantially accelerates optimization and improves the final fit compared to training the same architecture from scratch.
Figure S.1: Effect of foundation-model pretraining on optimization. Extending and reusing a pretrained MACE model on the curated dataset (blue curves) significantly accelerates convergence compared to training from scratch (orange curves). Both the classification loss and the score-matching loss start from lower initial values and decrease more rapidly, reaching smaller final values within fewer epochs.
Appendix B Effect of elastic-strain augmentation on prototype classification
To assess the role of elastic-strain augmentation, we retrained the $\log\!P$ model with the same architecture and hyperparameters but without the random elastic-strain transformations applied during pretraining and fine-tuning. We then evaluated both models on a uniaxial shock-compression trajectory of HCP Ti (the same trajectory as in the main text). As shown in Supplementary Fig. S.2, the model trained without elastic-strain augmentation systematically misclassifies the uniaxially compressed HCP region as the rhombohedral A_hR3_166 prototype (space group 166), effectively explaining the strain-induced distortions by switching to a different prototype rather than recognizing them as elastically deformed HCP. In contrast, the model trained with elastic-strain augmentation correctly preserves the HCP label throughout the shocked region for all frames. This ablation confirms that elastic-strain augmentation is critical for robust generalization to strongly compressed microstructures and suppresses spurious prototype switching under large uniaxial strain.
Figure S.2: Effect of elastic-strain augmentation on prototype classification in shocked HCP Ti. (a–c) Frames 0, 7, and 14 from a uniaxial shock-compression trajectory of HCP Ti (shock applied from the bottom), using a model retrained without random elastic-strain augmentation. The uniaxially compressed region is systematically misclassified as the rhombohedral A_hR3_166 prototype (space group 166, yellow), indicating that the network explains strain-induced distortions by switching prototypes rather than recognizing them as deformed HCP. (d–f) The same frames from the same trajectory evaluated with a model trained with elastic-strain augmentation, which correctly classifies the entire shocked region as HCP (red) and eliminates the spurious A_hR3_166 pocket, demonstrating that elastic-strain augmentation is essential for robust generalization to strongly compressed microstructures.